# AI-accelerated Discovery of Altermagnetic Materials

### v3

@@ Methods
The objective of our proposed pre-training model for crystal materials is to learn the information embedded within crystal structures. To enhance the learning capacity of the pre-training model, we proposed a graph auto-encoder architecture (see Fig. 1 and Fig. 2). The encoder consists of $n$ layers of graph convolution to learn crystal embeddings, while the decoder employs the Wasserstein distance based on the optimal transport theory [33] for the reconstruction of the input crystal structures. Specifically, the encoder aims to encode the graphical representation of crystal materials into a high-dimensional matrix, while the goal of the decoder is to decode this one back into the graphical representation of crystal materials. Through extensive training with unlabeled data, the model effectively converges (as depicted in the pre-training loss history as shown in Supplementary Fig. S.3). We believe that the pre-trained encoder can effectively project the crystal structures into crystal embeddings. Leveraging the encoder of the pre-trained model, we built the classifier model by incorporating a pooling layer and a softmax function. Subsequently, we trained the classifier model using the fine-tuning dataset. This trained model is then employed to screen the candidate materials, offering the probability of whether the target material is altermagnetic. The hyperparameters of the model were obtained by grid search, as listed in Supplementary Table S.1. In summary, our model comprises four main components: crystal data pre-processing, an encoder constructed using graph convolutional neural networks, a decoder built based on optimal transport theory, and the construction of a classifier model. We elaborate on each of these components one by one as follows.
Crystal data pre-processing.
@@
The data pre-processing procedure aims to bridge the crystal structure and the crystal graph representation [30]. The input of the model is a crystal structure file (.cif) that contains three primitive translation vectors of the primitive unit cell and the positional information of each atom. It satisfies the organization invariance for atomic indexing and the size invariance for unit cell selection. We define the graph representation $\mathcal{G}(V,U,X)$ to describe the crystal structure information, where $V$ denotes the set of nodes, $U$ the set of edges, and $X$ the set of features. First, we represent atoms as nodes $v_{i}$ in a crystal graph representation, where $i=1,...,|V|$. Since periodic boundary conditions are taken into consideration, equivalent nodes are merged to obtain irreducible nodes. Then, for each node $v_{i}$, we consider the neighborhood nodes $v_{j}$, where $j=,1,...,|\mathcal{N}_{i}|$ and $\mathcal{N}_{i}$ the set of neighborhood nodes for $v_{i}$. The $k$ connections between nodes $v_{i}$ and $v_{j}$ are denoted as the edge $u_{(i,j)_{k}}$ in the graph. Next, the initial node features $\big{\{}h_{i}^{(0)}\big{\}}_{i=1}^{|V|}$ are given through one-hot encoding based on the sequence of atoms in the crystal structure. We use $H_{\mathcal{N}_{i}}^{(0)}$ to denote the neighbor node features of node $v_{i}$. Here, $v_{i}$ denotes the $i$-th node. Each edge $u_{(i,j)_{k}}\in U$ is represented by a feature vector $u_{(i,j)_{k}}$ that corresponds to the $k$-th bond linking node $v_{i}$ and node $v_{j}$. A feature vector $h_{i}\in X$ encoding the attribute of the atom corresponding to node $v_{i}\in V$ is used to represent each node $v_{i}$. An example for determining the atom connectivity is illustrated in Supplementary Fig. S.4.
Crystal graph convolutional encoder.
@@
$h_{i}^{(t+1)}=h_{i}^{(t)}+\sum_{v_{j}\in\mathcal{N}_{i},v_{m}\in\mathcal{M}_{i},k}\sigma\left(z^{(t)}_{(i,j)_{k}}W_{f}^{(t)}+b_{f}^{(t)}\right)\odot g\left(z^{(t)}_{(i,j)_{k}}W_{s}^{(t)}+h_{i,m}^{(t)}W_{m}^{(t)}+b_{s}^{(t)}\right),$ (1)
where $\odot$ denotes the element-wise multiplication, $\mathcal{M}_{i}$ the magnetic atoms corresponding to node $v_{i}$, and $\sigma$ the sigmoid activation function. Since the magnetic atoms are important for the material to exhibit altermagetic properties, we added the weight term $W_{m}^{(t)}$ for the magnetic atoms. The weight function $W_{c}^{(t)},W_{s}^{(t)},W_{m}^{(t)}$ are the convolution weight matrix, self-weight matrix, and magnetic atom weight matrix of the $t$-th layer, respectively. In Eq. (1), we incorporate the residual term $h_{i}^{(t)}$ to enhance the training of the neural network.
Neighborhood Wasserstein reconstruction decoder.
@@
The pre-training model.
To extract the crystal embeddings of the candidate materials, we employ a graph convolution neural network as an encoder, which consists of 3 graph convolution layers [^2] We refer the code of CGCNN [30] to construct the encoder module, whose github link is https://github.com/txie-93/cgcnn and commit ID is “f42ab233c4ee0c416879d6bc2d22a264418413ad”.. For classification, we utilize a pooling layer and a multilayer perceptron as the projection head, comprising two fully connected layers with a ReLU activation layer and a Dropout layer [^3] We refer the code of NWR-GAE [39] to construct the auto-decoder module and compute loss, whose github link is https://github.com/mtang724/NWR-GAE and commit ID is “e9aee57a009f6c2150cf68fc173c2af3094a7205”. In terms of optimization, we use the Adam optimizer with a learning rate of 0.001. Additionally, we implement MultiStepLR with milestones set at 100. To ensure stability during training, we train our classifier model with a dropout rate of 0.25. This is done using a batch size of 512 and training over 500 epochs on 2 NVIDIA A100 GPUs. We then label these materials whose probabilities exceed 0.9 as potential altermagnetic candidates in all experiments. To train and evaluate our model efficiently, we leverage the distributed deep learning framework Accelerate. More details of the model training are discussed in Supplementary Note A. Given this specific network architecture, we can complete the training and evaluation process of the classifier model in less than 1.5 hours on our datasets. This significantly improves the overall efficiency of our workflow.
The first-principles electronic calculation.
The first-principles electronic structure calculations were performed in the framework of density functional theory (DFT) using the Vienna Abinitio Simulation Package (VASP) [49]. The generalized gradient approximation (GGA) of the Perdew-Burke-Ernzerhof (PBE) type was adopted for the exchange-correlation functional [50]. The projector augmented wave (PAW) method was employed to describe the interactions between valence electrons and nuclei [51]. To account for the correlation effects of 3d and 4f orbitals, we performed $\rm GGA+U$ calculations by using the simplified rotationally invariant version introduced by Dudarev et al. [41]. The detail information for DFT calculations is given in Supplement Note D.
Data availability
The crystal data are available from the Materials Project database via the web interface at https://materialsproject.org or the API at https://api.materialsproject.org.
Code availability

### v4

@@ Methods
The objective of our proposed pre-training model for crystal materials is to learn the information embedded within crystal structures. To enhance the learning capacity of the pre-training model, we proposed a graph auto-encoder architecture (see Fig. 1 and Fig. 2). The encoder consists of $n$ layers of graph convolution to learn crystal embeddings, while the decoder employs the Wasserstein distance based on the optimal transport theory [39] for the reconstruction of the input crystal structures. Specifically, the encoder aims to encode the graphical representation of crystal materials into a high-dimensional matrix, while the goal of the decoder is to decode this one back into the graphical representation of crystal materials. Through extensive training with unlabeled data, the model effectively converges (as depicted in the pre-training loss history as shown in Supplementary Appendix Fig. S.3). We believe that the pre-trained encoder can effectively project the crystal structures into crystal embeddings. Leveraging the encoder of the pre-trained model, we built the classifier model by incorporating a pooling layer and a softmax function. Subsequently, we trained the classifier model using the fine-tuning dataset. This trained model is then employed to screen the candidate materials, offering the probability of whether the target material is altermagnetic. The hyperparameters of the model were obtained by grid search, as listed in Supplementary Appendix Table S.1. In summary, our model comprises four main components: crystal data pre-processing, an encoder constructed using graph convolutional neural networks, a decoder built based on optimal transport theory, and the construction of a classifier model. We elaborate on each of these components one by one as follows.
Crystal data pre-processing
@@
The data pre-processing procedure aims to bridge the crystal structure and the crystal graph representation [36]. The input of the model is a crystal structure file (.cif) that contains three primitive translation vectors of the primitive unit cell and the positional information of each atom. It satisfies the organization invariance for atomic indexing and the size invariance for unit cell selection. We define the graph representation $\mathcal{G}(V,U,X)$ to describe the crystal structure information, where $V$ denotes the set of nodes, $U$ the set of edges, and $X$ the set of features. First, we represent atoms as nodes $v_{i}$ in a crystal graph representation, where $i=1,...,|V|$. Since periodic boundary conditions are taken into consideration, equivalent nodes are merged to obtain irreducible nodes. Then, for each node $v_{i}$, we consider the neighborhood nodes $v_{j}$, where $j=,1,...,|\mathcal{N}_{i}|$ and $\mathcal{N}_{i}$ the set of neighborhood nodes for $v_{i}$. The $k$ connections between nodes $v_{i}$ and $v_{j}$ are denoted as the edge $u_{(i,j)_{k}}$ in the graph. Next, the initial node features $\big{\{}h_{i}^{(0)}\big{\}}_{i=1}^{|V|}$ are given through one-hot encoding based on the sequence of atoms in the crystal structure. We use $H_{\mathcal{N}_{i}}^{(0)}$ to denote the neighbor node features of node $v_{i}$. Here, $v_{i}$ denotes the $i$-th node. Each edge $u_{(i,j)_{k}}\in U$ is represented by a feature vector $u_{(i,j)_{k}}$ that corresponds to the $k$-th bond linking node $v_{i}$ and node $v_{j}$. A feature vector $h_{i}\in X$ encoding the attribute of the atom corresponding to node $v_{i}\in V$ is used to represent each node $v_{i}$. An example for determining the atom connectivity is illustrated in Supplementary Appendix Fig. S.4.
Crystal graph convolutional encoder
@@
$h_{i}^{(t+1)}=h_{i}^{(t)}+\sum_{v_{j}\in\mathcal{N}_{i},v_{m}\in\mathcal{M}_{i},k}\sigma\left(A\right)\odot g\left(B\right),$ (1)
where $A$ denotes $z^{(t)}_{(i,j)_{k}}W_{f}^{(t)}+b_{f}^{(t)}$ and $B$ denotes the $z^{(t)}_{(i,j)_{k}}W_{s}^{(t)}+h_{i,m}^{(t)}W_{m}^{(t)}+b_{s}^{(t)}$. The $\odot$ denotes the element-wise multiplication, $\mathcal{M}_{i}$ the magnetic atoms corresponding to node $v_{i}$, and $\sigma$ the sigmoid activation function. Since the magnetic atoms are important for the material to exhibit altermagetic properties, we added the weight term $W_{m}^{(t)}$ for the magnetic atoms. The weight function $W_{c}^{(t)},W_{s}^{(t)},W_{m}^{(t)}$ are the convolution weight matrix, self-weight matrix, and magnetic atom weight matrix of the $t$-th layer, respectively. In Eq. (1), we incorporate the residual term $h_{i}^{(t)}$ to enhance the training of the neural network.
Neighborhood Wasserstein reconstruction decoder
@@
The pre-training model
To extract the crystal embeddings of the candidate materials, we employ a graph convolution neural network as an encoder, which consists of 3 graph convolution layers [^5] We refer the code of CGCNN [36] to construct the encoder module, whose github link is https://github.com/txie-93/cgcnn and commit ID is “f42ab233c4ee0c416879d6bc2d22a264418413ad”.. For classification, we utilize a pooling layer and a multilayer perceptron as the projection head, comprising two fully connected layers with a ReLU activation layer and a Dropout layer [^6] We refer the code of NWR-GAE [49] to construct the auto-decoder module and compute loss, whose github link is https://github.com/mtang724/NWR-GAE and commit ID is “e9aee57a009f6c2150cf68fc173c2af3094a7205”. In terms of optimization, we use the Adam optimizer with a learning rate of 0.001. Additionally, we implement MultiStepLR with milestones set at 100. To ensure stability during training, we train our classifier model with a dropout rate of 0.25. This is done using a batch size of 512 and training over 500 epochs on 2 NVIDIA A100 GPUs. We then label these materials whose probabilities exceed 0.9 as potential altermagnetic candidates in all experiments. To train and evaluate our model efficiently, we leverage the distributed deep learning framework Accelerate. More details of the model training are discussed in Supplementary Appendix Note A. Given this specific network architecture, we can complete the training and evaluation process of the classifier model in less than 1.5 hours on our datasets. This significantly improves the overall efficiency of our proposed workflow.
The first-principles electronic calculation
The first-principles electronic structure calculations were performed in the framework of density functional theory (DFT) using the Vienna Abinitio Simulation Package (VASP) [92]. The generalized gradient approximation (GGA) of the Perdew-Burke-Ernzerhof (PBE) type was adopted for the exchange-correlation functional [93]. The projector augmented wave (PAW) method was employed to describe the interactions between valence electrons and nuclei [94]. To account for the correlation effects of 3d and 4f orbitals, we performed $\rm GGA+U$ calculations by using the simplified rotationally invariant version introduced by Dudarev et al. [84]. The detail information for DFT calculations is given in Supplement Appendix Note D.
SUPPLEMENTARY DATA
Supplementary data are available at NSR online. The crystal data are available from the Materials Project database via the web interface at https://materialsproject.org or the API at https://api.materialsproject.org.